Jane has some nickels and quarters. If she has 3 more nickels than she has quarters and the value of the coins is $4.05 how many nickels and quarters does Jane have?
Answer by stanbon(75887)

Jane has some nickels and quarters. If she has 3 more nickels than she has quarters and the value of the coins is $4.05 how many nickels and quarters does Jane have?
-----------------------------------
Quantity Equation: q = n + 3
Value Equation:::25q + 5n = 405 cents
-----
Multiply thru the Quantity Eq. by 5 to get:
5q-5n= 15
----------------
Add that to the Value Eq:
30q = 420
q = 14 (# of quarters)
----------------------------
Substitute into the Quantity Eq to solve for "n":
14 = n + 3
n = 11 (# of nickels)
============================
Cheers,
Stan H.